Ryan invested some money at 9% and twice as much at 10%. The total income for investment every year is 5 220 pesos. How much total money did he have in the beginning?


Pls. ans. algebraically..

Answer by amoresroy(361)

Ryan invested some money at 9% and twice as much at 10%. The total income for investment every year is 5 220 pesos. How much total money did he have in the beginning?
Let x = the money invested at 9%
2x = the money invested at 10%
x+2x = the total money in the beginning
.09x + .10(2x) = 5220
.29x = 5220
x = 5220/.29
x = 18,000
x+2x = 18,000 + 2(18,000) = 54,000
He had P54,000 in the beginning.